OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

(5 + √3) Part 1: (5 - √3) This is the conjugate because in a binomial, the conjugate has the same terms, but a different sign in between them. Part 2: (5 + √3)(5 - √3) = 25 - 3 = 22